Five hundred children participated in a field demonstration. Their heights averaged is 110 cm with a standard deviation of 6 cm. What is the probability that the height of a child, picked at random, is greater than 104 cm?


1
Expert's answer
2021-03-26T03:18:44-0400

Here the everage heights of the "500" children is "110" cm with standard deviation "6" cm.

Let "X" be a random variable denotes the height of the students.

Then "X" is normally distributed with mean "110" cm and standard deviation "6" cm.

Therefore we have "\mu =110" and "\sigma=6" .

Let us take "Z= \frac{X-\mu}{\sigma}" . Then "Z=\frac{X-110}{6}".

Now we have to find "P(X>104)."

"\therefore P(X>104)=P(Z>\frac{104-110}{6})"

"=P(Z>-1)"

"=0.5+P(0\leq Z\leq 1)"

"=0.5+0.3413" [ from normal distribution table ]

"=0.8413"

Which is the required probability.